HETAV PATEL

Grade 11,

if the time period of revolution of a planet is increased to 3 root3 times its present value then from what percentage its radius of the orbit of revolution will increases

if the time period of revolution of a planet is increased to 3 root3 times its present value then from what percentage its radius of the orbit of revolution will increases

Grade:9

1 Answers

shashwat giri
6 years ago
as we know the third kepler’s law of planetry motion that is T2 = R3
so according to that 
T=   K * R3       [1]
{3\sqrt{}3}2 =  K*{3\sqrt{}3}3
ATQ!
{3\sqrt{}3}3/t3 *100
 
T = 100\sqrt{}3